Angie’s Bakery bakes bagels. The June production is given below.

250 370 370 250 250
210 370 370 370 250
250 370 250 370 370
370 370 370 180 180
210 370 210 230 180
180 370 370 370 180


Find the weighted mean. (Round answer to the nearest whole number.)


Weighted mean             bagels
Answer #1

Answer : 294

To find the weighted mean, we first find the frequency of each value. The frequency is as below

Production Frequency
250 6
210 3
370 15
180 5
230 1

We then find the sum of product of all the production amount and the frequency. This is equal to below

Sumproduct = 250*6+210*3+370*15+180*5+230*1 = 1500+630+5550+900+230 = 8810.

To find the weighted mean, we then divide this sum by the total of frequency = 8810/30 = 293.66 = 294